OpenStudy (anonymous):

how to solve this quadratic 143 = (9+2x)(7+2x)

OpenStudy (anonymous):

-4x^2-32x+80=0 x=-10 x=2

OpenStudy (anonymous):

How did you get negatives when everything in the parenthesis is positive?

OpenStudy (anonymous):

first multiply (9+2x) and (7+2x) then subtract 143 from both sides u will get the quadratic equation in standard form, then solve it with quadratic formula.

OpenStudy (campbell_st):

expand the right hand side 143 = 4x^2 +32x + 63 subtract 143 from both sides 4x^2 + 32x - 80 = 0 factorising 4(x^2 + 8x - 20) = 0 this factorises to 4(x+10)(x -2)=0 x = -10, 2